OpenStudy (anonymous):

In the class elections, Connie received 53 more votes than Carl. In all, 221 votes were cast for the two candidates. How many votes did each receive. In the class elections, Connie received 53 more votes than Carl. In all, 221 votes were cast for the two candidates. How many votes did each receive. @Mathematics

OpenStudy (anonymous):

CarlVotes+(ConnieVotes+53)=221 x+(x+53)=221 2x=221-53 2x=168 x= 84 vote for carl add the 53 votes to the 84 to get Connies votes 53 + 84 = 137 to check ( 137(connies) + 84(carls)= 221 total votes )